# Bitangent surfaces and involutions of quartic surfaces

### v2

@@ 7. Kummer quartic surfaces in $p=2$: $2$-rank $1$ case
where $\alpha,\beta$ are constants with $\beta\neq 0$ (Duquesne [7, §3]). The surface $X_{1}$ has exactly two singular points
@@
of type $D_{8}$ (Shioda [21], Katsura [12]) and contains two tropes cutting by the hyperplane section $x=0$ and $z=0$, respectively. Two tropes meet at $p_{1}$ and $p_{2}$.
@@
Thus, the congruence of lines $S(\sigma_{2})$ is a quadric cone, a special linear section of the Grassmannan quadric.
@@
It is enough to show that the class $n=4$, that is, a general hyperplane section of $X_{1}$ is a plane quartic with exactly four bitangent lines. This follows from the result of Stöhr-Voloch [22, p.60]. We can calculate the Hasse-Witt invariant of the hyperplane section which is equal to 2. Then, it follows from [22, p.60] that the number of bitangent lines of the hyperplane section is equal to 4. ∎
@@
Let $\ell_{0}=V(x,z)$ be the line which is the intersection of two trope-hyperplanes. Then, any bitangent line defined by $\sigma_{2}$ meets $\ell_{0}$ as follows. Let $p=[x_{0},y_{0},z_{0},w_{0}]\in\mathbb{P}^{3}$ be a general point and $q=\sigma_{2}(p)$. The bitangent line $\langle p,q\rangle$ is given by
@@
8. Kummer quartic surfaces in $p=2$: supersingular case
@@
It presereves $X_{0}$, and restricts to a bitangent involution $\sigma$ of $X_{0}$. This follows from a direct calculation.
@@
Let $X_{0}$ be the Kummer quartic surface associated with a supersingular curve. Then, ${\mathrm{Bit}}(X_{0})$ is of bidegree $(1,2)$ in $G_{1}(\mathbb{P}^{3})$. It consists of a trope and a quadric cone.
@@
It is enough to show that the class $n=2$, that is, a general hyperplane section of $X_{0}$ is a plane quartic with exactly two bitangent lines. This follows from the result of Stöhr-Voloch [22, p.60], too. We can calculate the Hasse-Witt invariant of the hyperplane section which is equal to 1. Then, it follows from [22, p.60] that the number of bitangent lines of the hyperplane section is equal to 2. ∎
@@
Theorem 8.3.
Let $X$ be a non-conical normal quartic surface $X$ over a field of characteristic $2$. Then, its general plane section is a plane quartic curve with positive $2$-rank of its Jacobian.
Proof.
A general plane section of $X$ is a smooth quartic curve. Suppose it is supersingular. A plane quartic curve over any field $K$ (not necessary algebraically closed) is isomorphic to a curve
$Q(x,y,z)^{2}+x(y^{3}+zx^{2})=0,$
where $Q$ is a quadratic form with a non-zero coefficient at $z^{2}$ [18, Proposition 2.4]. Its unique bitangent line is $V(x)$.
[…]

### v3

@@ 7. Kummer quartic surfaces in $p=2$, $2$-rank $1$ case
where $\alpha,\beta$ are constants with $\beta\neq 0$ [7, §3]. The surface $X_{1}$ has exactly two singular points
@@
of type $D_{8}$ [21], [12] and contains two tropes defined by the hyperplane section $x=0$ and $z=0$, respectively. The two tropes meet at $p_{1}$ and $p_{2}$.
@@
Thus, the congruence of lines $S(\sigma_{2})$ is a quadric cone, a special linear section of the Grassmannian quadric.
@@
We know that ${\mathrm{Bit}}(X)$ contains two $\beta$-planes and two irreducible components of bidegree $(1,1)$. Taking the partial derivatives, we find that they are linearly independent, and hence, there are no inseparable projection centers. So, ${\mathrm{Bit}}(X)$ does not contain $\alpha$-planes. The equation of $X$ is of the form
$Q^{2}+x^{2}zw+xyz^{2}=0.$
Substituting $w=ax+by+cz$, we obtain the equation of a general plane section $q^{2}+x^{2}z(ax+by)+xyz^{2}=0$. It is immediate to check that it is projectively equivalent to a plane quartic of type (2) from Proposition 5.8. Thus, its $2$-rank is equal to $4$, and hence, the class of ${\mathrm{Bit}}(X)$ is equal to $4$. So, we have found all irreducible components of ${\mathrm{Bit}}(X)$: two $\beta$-planes and two components of bidegree $(1,1)$. ∎
@@
Let $\ell_{0}=V(x,z)$ be the line which is the intersection of the two trope-hyperplanes. Then, any bitangent line defined by $\sigma_{2}$ meets $\ell_{0}$ as follows. Let $p=[x_{0},y_{0},z_{0},w_{0}]\in\mathbb{P}^{3}$ be a general point and $q=\sigma_{2}(p)$. The bitangent line $\langle p,q\rangle$ is given by
@@
8. Kummer quartic surfaces in $p=2$, supersingular case
@@
The involution $T$ preserves $X_{0}$, and restricts to a bitangent involution $\sigma$ of $X_{0}$. This follows from a direct calculation.
@@
Let $X_{0}$ be the Kummer quartic surface associated with a supersingular curve. Then, ${\mathrm{Bit}}(X_{0})$ is of bidegree $(1,2)$ in $G_{1}(\mathbb{P}^{3})$ and it consists of a trope and a quadric cone.
@@
Taking the partial derivatives, we find again that there are no inseparable projection centers. So, ${\mathrm{Bit}}(X)$ does not contain $\alpha$-planes. Since we have already found irreducible components of ${\mathrm{Bit}}(X)$ of bidegree $(0,1)$ and $(1,1)$, it is enough to check that the $2$-rank of a general plane section is equal to $2$. This can be seen directly by plugging in $w=ax+by+cz$ in the equation of $X$ and reducing the non-square part of the obtained equation to the form (3) from Proposition 5.8. ∎